3.4 Assess Your Understanding
Applying the Concepts #5
Applying the Concepts NW 5. Birth Weights Babies born after a gestation period of 32-35 weeks have a mean weight of 2600 grams and a standard deviation of 660 grams. Babies born after a gestation period of 40 weeks have a mean weight of 3500 grams and a standard deviation of 470 grams. Suppose a 34-week gestation period baby weighs 2400 grams and a 40-week gestation period baby weighs 3300 grams. What is the z-score for the 34-week gestation period baby? What is the z-score for the 40-week gestation period baby? Which baby weighs less relative to the gestation period?
Answer #1

Answer :

given data :-

mean weight = 3500 grams

mean weight = 3500 grams

mean weight = 3500 grams

standard deviation = 660

babies born after a period of 32-35 weeks

Babies born after gestation period of 40 weeks

have an mean weight = 3500 grams

\mu =3500

mean weight = 3500 grams

standard deviation = 470 grams

  \sigma = 470

a 40 week gestation period baby weighs = 3300 grams

Z score for X =3300 grams

we know that

Z = X- \mu/ \sigma

= 3300-3500/470

= -200/470

Z = -0.4255

34 week gestation period baby weighs

X = 2400 grams

have an mean weight = 3500 grams

standard deviation = 660

Z = X- \mu/ \sigma

Z= 2400-2600/660

Z= -200/660

Z= -0.3030

\therefore THE 40 WEEKS BABY WEIGHS LESS RELATIVE TO GESTATION PERIOD.